OpenStudy (anonymous):

how long will it take $5000 to double if it is invested at 6% interest compounded continually? (A=Pe^rt)

OpenStudy (wolf1728):

The formula is years = log (total) - log(principal) / log(1+rate) First, let's figure out 6% compounded continuously Annual Rate = e^r -1 Annual Rate = (2.718281828^.06)-1 Annual Rate = 0.0618365465 years = [log (10,000) - log (5,000)] / log(1.0618365465) years = (4 - 3.6989700043) / 0.0260576689 years = 0.3010299957 / 0.0260576689 years = 11.5525